The perimeter of a standard high school basketball court is 268 ft. The length is 34 ft greater than the width. Find the length and the width of the court.
Answer by nerdybill(7384)

The perimeter of a standard high school basketball court is 268 ft. The length is 34 ft greater than the width. Find the length and the width of the court.
.
Let w = width
then
w+34 = length
.
2(w + w+34) = 268
dividing both sides by 2:
(w + w+34) = 134
2w+34 = 134
2w = 100
w = 100 ft. (width
.
Length:
w+34 = 100+34 = 134 ft